ROTATIONAL SPECTRA AND STRUCTURE OF H_{2}S COMPLEXES: Ar(H_{2}S)_{2} AND C_{2}H_{4}-H_{2}S

\underline{E. ARUNAN}, P. K. MANDAL, MAUSUMI GOSWAMI, Department of Inorganic and Physical Chemistry, Indian Institute of Science, Bangalore 560 012 INDIA.

Recently, we determined hydrogen bond radii for HF, HCl, HBr, HI, HCN, H_{2}O and HCCH, using the accurate structural data available from microwave spectroscopic studies on several B:HX complexes, where B is a hydrogen bond acceptor^{a}. For H_{2}S, the available data is limited compared to other hydrogen bond donors listed above. Hence, a systematic investigation has been started on several H_{2}S complexes. Several transitions have been observed for Ar-(H_{2}S)_{2} and C_{2}H_{4}- H_{2}S complexes. The rotational spectrum of C_{2}H_{4}-H_{2}S complex appears to be a composite of rotational spectra of the hydrogen bonded C_{2}H_{4}-H_{2}O and the van der Waals complex C_{2}H_{4}-Ar. Each transition is split in to four with a smaller splitting of about 0.14 MHz and a larger splitting of 1.67 MHz in (B+C)/2. The rotational constants for the strongest progression are A = 26(1) GHz; B = 1972.88(1) MHz and c = 1866.68(1) MHz. Assignment of Ar-(H_{2}S)_{2} spectra is in progress. It was realized that the rotational spectrum for (H_{2}S)_{2} complex has not been completely solved yet^{b} and hence a ’revisit’ started. Several more progressions have been assigned for this dimer now. The search for deuterated isotopomers was carried out by bubbling H_{2}S through D_{2}O and some new H_{2}O-H_{2}S transitions have been observed. During this talk, rotational spectra of all these complexes will be discussed.

aP. K. Mandal and E. Arunan J. Chem. Phys. 114, 3880 (2001)

bF. J. Lovas private communication